A garden alongside the wall of a house is to be fenced in using 100 feet of fencing. If the wall of the house is one side o the garden, determine the greatest possible area that can be enclosed
Answer by checkley77(12569)

DIVIDE THE FENCING BY 4.
100/4=25. THIS WILL BE THE WIDTH OF THE 2 ENDS.
THE LENGTH WILL BE DOUBLE THE WIDTH.
2L=2*25=50 ANS. IS THE LENGTH.
LW=50*25=1,250 FT^2.
PROOF:
REDUCE THE WIDTH & INCREASE THE LENGTH.
24*52=1,248 FT^2.
REDUCE THE LENGTH & INCREASE THE WIDTH
26*48=1,248 FT^2.